RhondaSommer (rhondasommer):

Using the data, determine a function in the form f ( x ) = A sin(k(x – c) ) + h . Amplitude: 26 Period: 12 Phase Shift: 56

RhondaSommer (rhondasommer):

@3mar

RhondaSommer (rhondasommer):

@zepdrix

OpenStudy (princeharryyy):

amplitude is given so u do not have to do any thing in that. ok.

OpenStudy (princeharryyy):

next, how to get a period of 12 right.

OpenStudy (princeharryyy):

Amplitude stays what has been given that is A =26

OpenStudy (princeharryyy):

next u need to find k before you do that. here's an example. let's say fuction is y = sin2x, it's period is 2pie/2 =pie generalizing the fact if u have any function say y= sinkx, it's period is always 2pie/k, u just divide it by the number so because you need to find k in your question.

OpenStudy (princeharryyy):

you are given period = 12 => 12 = 2pie/k => k =6/pie

OpenStudy (princeharryyy):

wait, k = pie/6 (correction, I did calculation wrong earlier) next you need to find c, so whenever you have any function for example y= sin(B(x-C)) the phase shift becomes C/B now as we have already find k, and you are given phase shift 56 using the above formula 56 = C/k => 56 = C/(pie/6) => 56*pie/6 =C => C = 56pie/6 = 28pie/3

OpenStudy (princeharryyy):

areyou getting it @RhondaSommer

OpenStudy (princeharryyy):